7.7 A certain compressible layer has a thickness of 4m. After 1 year when the clay is 50% consolidated, 8 cm of settlement has occurred. For similar clay and loading conditions, how much settlement would occur at the end of 1 year and 4 years if the thickness of this new layer were 40 m?
Answer #1

Solution:

we know that loading condition and clay condition are similar so we can say that

4m thickness of clay layer settle in one year = 8/100 m

therefore,

1m thickness of clay layer settle in one year = 8/(100x4)

so,

40m thickness of clay layer settle in one year = (8/(100 x 4) )x 40 = .80m or 80 cm

for 4 year the settlement becomes 4 times of one year settlement = 80x4 = 320 cm